# Adaptive Parametric Activation

### v2

@@ 0.C.3 Impact of Initialisation
$Uniform(-2,0)$ 57.3
$Uniform(-3,0)$ 57.4
$Uniform(-2,2)$ Failed
$Uniform(0.8,1.2)$ 57.9
Table 12: $\kappa$ parameter initialisation inside the attention layer, using APA* ResNet50 backbone on ImageNet-LT. The $\lambda$ is initialised with $Uniform(0,1)$ by default.
@@
$Uniform(-1,0)$ 57.9
0.C.4 Channel specific $\lambda$ and $\kappa$
We have also tried a variant that uses seperate $\lambda$ and $\kappa$ parameters for every channel. As shown in Table 13, this variant performs worse than using shared $\lambda$ and $\kappa$ parameters for the channels.
Table 13: Results with Channel Specific $\lambda$ and $\kappa$, using APA* ResNet50 backbone on ImageNet-LT.
@@ 0.C.4 Channel specific $\lambda$ and $\kappa$
0.C.5 Baseline enhancements
We show the detailed ablation study for ImageNet-LT in Table 14. First, the vanilla ResNet50 model trained for 100 epochs on ImageNet-LT achieves $44.4\%$. When we train for 200 epochs then it adds $1.5$pp and switching from linear classifier to cosine classifier adds another $0.4$pp. Stronger training techniques such Mixup [106], Auto-Augment [12] and weight decay tuning further boost the performance by $3.3$pp. Post-calibrated Softmax [34] adds an additional $5.4$pp and finally the Squeeze and Excite module [37] adds another $1.0$pp reaching the final $56.0\%$. Most baseline performance comes from the PC-Softmax and the weight decay finetuning. On top of this strong baseline, our APA increases the performance by $1.0$pp, showing its strong generalisability. Dropout and LayerNorm further increase the performance by $0.4$pp and finally AGLU adds a respectable $0.5$pp reaching $57.9\%$ accuracy on ImageNet-LT. The absolute improvement of all modules is $13.5$pp and our proposed methods, APA and AGLU, contribute by $1.5$pp which is a relative $11\%$ of the total absolute improvement.
Table 14: Detailed Ablation Study, using ResNet50 on ImageNet-LT.
200 epochs Cosine Classifier SE-nets[37] AutoAugment[12] Mixup [106] Weight Decay Tuning [3] PCS [34] APA (ours) Dropout [32] LayerNorm [4] AGLU (ours) ImagetNet-LT
@@ 0.C.6 Qualitative Results
0.C.6.1 Visualisations on Imagenet-LT
We further show more qualitative results on ImageNet-LT with ResNet50 backbone. On the left subfigure, we show the model’s highest prediction marked with F,C,R that stands for frequent, common and rare class respectively and the Grad-cam activation [75]. On the right subfigure, we show the last layer’s channel attention signal and its corresponding entropy denoted with (E). As the Figure shows, APA* produces higher entropy attention signals than the baseline and predicts both frequent and rare classes correctly.
@@ 0.C.7 Calibration results
0.C.8 Next textual token prediction experiment
We perform one preliminary next-token prediction experiment using GPT2 [70] and the FineWeb-Edu [59] subset that contains 10 billion GPT2 tokens. The model is based on the GPT2 smallest architecture, which contains 117M parameters, and the code implementation follows [45]. We train the GPT2 model for one epoch, using 8 V100 GPUs, a total batch size of 0.5M tokens, learning rate $46e-$ 4 and Adam optimizer [48] with momentum. We test the model on the HellaSwag benchmark [105] using zero-shot evaluation. To apply AGLU with GPT2, we simply switch the GELU activation with AGLU inside all MLP layers of the transformer. As the results suggest in Table 15, our AGLU increases the performance of GPT2 by $0.4\%$, showing that AGLU could be a good alternative for text-classification.
Table 15: Comparative results using GPT2 smallest model and HellaSwag benchmark.
Method Accuracy
GELU 31.0
AGLU 31.4

### v3

@@ C-A Impact of Initialisation
Uniform(2,0)Uniform(-2,0) 57.3
Uniform(3,0)Uniform(-3,0) 57.4
Uniform(2,2)Uniform(-2,2) Failed
$Uniform(1,1.3)$ 57.9
TABLE XVII: $\kappa$ parameter initialisation inside the attention layer, using APA* ResNet50 backbone on ImageNet-LT. The $\lambda$ is initialised with $Uniform(0,1)$ by default.
@@
Uniform(1,0)Uniform(-1,0) 57.9
C-B Channel specific $\lambda$ and $\kappa$
We have also tried a variant that uses seperate $\lambda$ and $\kappa$ parameters for every channel. As shown in Table XVIII, this variant performs worse than using shared $\lambda$ and $\kappa$ parameters for the channels.
TABLE XVIII: Results with Channel Specific $\lambda$ and $\kappa$, using APA* ResNet50 backbone on ImageNet-LT.
@@ C-B Channel specific $\lambda$ and $\kappa$
C-C Baseline enhancements
We show the detailed ablation study for ImageNet-LT in Table XIX. First, the vanilla ResNet50 model trained for 100 epochs on ImageNet-LT achieves $44.4\%$. When we train for 200 epochs then it adds 1.51.5pp and switching from linear classifier to cosine classifier adds another 0.40.4pp. Stronger training techniques such Mixup [zhang2017mixup], Auto-Augment [cubuk2019autoaugment] and weight decay tuning further boost the performance by 3.33.3pp. Post-calibrated Softmax [hong2021disentangling] adds an additional 5.45.4pp and finally the Squeeze and Excite module [hu2018squeeze] adds another 1.01.0pp reaching the final $56.0\%$. Most baseline performance comes from the PC-Softmax and the weight decay finetuning. On top of this strong baseline, our APA increases the performance by 1.01.0pp, showing its strong generalisability. Dropout and LayerNorm further increase the performance by 0.40.4pp and finally AGLU adds a respectable 0.50.5pp reaching $57.9\%$ accuracy on ImageNet-LT. The absolute improvement of all modules is 13.513.5pp and our proposed methods, APA and AGLU, contribute by 1.51.5pp which is a relative $11\%$ of the total absolute improvement.
TABLE XIX: Detailed Ablation Study, using ResNet50 on ImageNet-LT.
200 epochs Cosine Classifier SE-nets[hu2018squeeze] AutoAugment[cubuk2019autoaugment] Mixup [zhang2017mixup] Weight Decay Tuning [alshammari2022long] PCS [hong2021disentangling] APA (ours) Dropout [hinton2012improving] LayerNorm [ba2016layer] AGLU (ours) ImagetNet-LT
@@ C-D Qualitative Results
C-D1 Visualisations on Imagenet-LT
We further show more qualitative results on ImageNet-LT with ResNet50 backbone. On the left subfigure, we show the model’s highest prediction marked with F,C,R that stands for frequent, common and rare class respectively and the Grad-cam activation [selvaraju2017grad]. On the right subfigure, we show the last layer’s channel attention signal and its corresponding entropy denoted with (E). As the Figure shows, APA* produces higher entropy attention signals than the baseline and predicts both frequent and rare classes correctly.